how can I write -2x^2+12x+12 in standard form?
Answer by khurshidi(7)

If the question is -2x^2+12x+12=0
Then you can take '-2' as common from the equation
i.e.
-2(x^2-6x-6)=0
-2 will go to the right hand side in the denominator of zero and will become zero,
so the equation will become
x^2 - 6x - 6 = 0
or x^2 + (-6)x + (-6) = 0
This is the standard form of quadratic equation.